Solve the inequality for: 8or=8x+37
Answer by Alan3354(69443)

Solve the inequality for: 8 <= 4-y
Add y
8 + y <= 4
Subtract 8
y <= -4
--------------
1-x >= 8x+37
Add x
1 >= 9x + 37
Subtract 37
-36 >= 9x
Divide by 9
-4 >= x
or x <= -4